OpenStudy (anonymous):

Difficult.... \[\sqrt{-147}=\sqrt{-147\times3\times-1}\]\[\sqrt{?}\times \sqrt{3}\times \sqrt{?}\]\[two blank spaces \sqrt{3}\] Help?

OpenStudy (anonymous):

something wrong here

OpenStudy (anonymous):

\[\sqrt{-147}=\sqrt{-147\times3\times-1}\] makes no sense

OpenStudy (angela210793):

sqr(-147)=isqr(49*3)=sqr48*sqr38sqr(-1)=7isqr3

OpenStudy (anonymous):

hmm where that negative 147 is under the square root with 3 was a blank space .

OpenStudy (anonymous):

oh i see you are supposed to recognize (how i don't know) that \[147=49\times 3\]

OpenStudy (anonymous):

so it should look like \[\sqrt{-147}=\sqrt{49\times 3\times -1}\]

OpenStudy (anonymous):

\[\sqrt{49}\times \sqrt{3}\times \sqrt{-1}=7 \sqrt{3}i\]